As a result of a ten percent increase in the price of pretzels, the weekly quantity demanded of potato chips increases from 700 bags to 900 bags. The cross price elasticity of demand for potato chips is:

2

2.5

.10

3.33

.4

1.82
Answer #1

Cross price Ed = % change in quantity demanded of potato chips / % change in the price of pretzels

% change in quantity demanded of potato chips = (900 - 700)/700 x 100 = 200/7 = 28.57

% change in price = 10

Cross price Ed = 28.57/10 = 2.857

Answer is close to 2.5